A buffer of Tris is made to a final pH of 7.5. At this pH, will the majority of Tris molecules be in the protonated or un-protonated form? The Ka of Tris-HCl (the protonated form) is 8.71 X10- 9 ?
Answer #1


Ka = 8.71*10^-9

pKa = - log (Ka)
= - log(8.71*10^-9)
= 8.06

use:
pH = pKa + log {[conjugate base]/[acid]}
7.5 = 8.06+log {[A-]/[HA]}
log {[A-]/[HA]} = -0.56
[A-]/[HA] = 0.2754
So,
[A-] < [HA]


Answer: Majority is in protonated form